OpenStudy (anonymous):

break this down using subtracting polynomials::: 3x(x-2)+x(4x-3)=

OpenStudy (anonymous):

first step is to multiply out using the distributive law. second step is to combine like terms

OpenStudy (anonymous):

\[3x(x-2)+x(4x-3)=3x^2-6x+4x^2-3x\]

OpenStudy (anonymous):

second step \[7x^2-9x\] steps clear?

OpenStudy (anonymous):

3x(x-2)=3x^2-6x x(4x-3)=4x^2-3x <<< add 7x^2-9x